# Linking numbers of modular knots

### 02 Jan 2024 v2

@@ 3. From the geometric cosine to the combinatorial cosign
The identity map for $q\in\mathbb{C}$ with $\lvert q\rvert<2$ and…
some graphs of $\operatorname{L}_{q}(A,B)$ for for $q\in\mathbb{C}$ with $\lvert q\rvert<1.6$.
@@ 2 $\operatorname{cosign}$ $\operatorname{cross}$ $\mathbf{+1}$ $\mathbf{0}$ $\mathbf{-1}$
For hyperbolic $A,B\in\operatorname{PSL}_{2}(\mathbb{Z})$ with axes $\alpha_{\mathcal{T}}=(\alpha_{-},\alpha_{+})$ and $\beta_{\mathcal{T}}=(\beta_{-},\beta_{+})$ in $\mathcal{T}$, we write $\operatorname{cross}(A,B)=\operatorname{cross}(\alpha_{+},\alpha_{-},\beta_{+},\beta_{-})$ and $\operatorname{cosign}(A,B)=\operatorname{cosign}(\alpha_{\mathcal{T}},\beta_{\mathcal{T}})$. Notice $\operatorname{cosign}(A,B)=1$ if and only if there exists $C\in\operatorname{PSL}_{2}(\mathbb{Z})$ such that $CAC^{-1},CBC^{-1}\in\operatorname{PSL}_{2}(\mathbb{N})$, in which case the set of such $C$ corresponds to the edges in $\alpha_{\mathcal{T}}\cap\beta_{\mathcal{T}}\subset\mathcal{T}$.
@@
The identity map for $q\in\mathbb{C}$ with $\lvert q\rvert<2$ and…
some graphs of $\operatorname{L}_{q}(A,B)$ for for $q\in\mathbb{C}$ with $\lvert q\rvert<1.6$.

### 04 Mar 2024 v4

@@ 3. From the geometric cosine to the combinatorial cosign
The identity map and some graphs of $\operatorname{L}_{q}(A,B)$ for $q\in\mathbb{C}$ with $\lvert q\rvert<1+\epsilon$.
@@ 2 $\operatorname{cosign}$ $\operatorname{cross}$ $\mathbf{+1}$ $\mathbf{0}$ $\mathbf{-1}$
For hyperbolic $A,B\in\operatorname{PSL}_{2}(\mathbb{Z})$ with axes $\alpha_{\mathcal{T}}=(\alpha_{-},\alpha_{+})$ and $\beta_{\mathcal{T}}=(\beta_{-},\beta_{+})$ in $\mathcal{T}$, we denote by $\operatorname{cross}(A,B)=\operatorname{cross}(\alpha_{+},\alpha_{-},\beta_{+},\beta_{-})$ and $\operatorname{cosign}(A,B)=\operatorname{cosign}(\alpha_{\mathcal{T}},\beta_{\mathcal{T}})$. Notice $\operatorname{cosign}(A,B)=1$ if and only if there exists $C\in\operatorname{PSL}_{2}(\mathbb{Z})$ such that $CAC^{-1},CBC^{-1}\in\operatorname{PSL}_{2}(\mathbb{N})$, in which case the set of such $C$ corresponds to the edges in $\alpha_{\mathcal{T}}\cap\beta_{\mathcal{T}}\subset\mathcal{T}$.
@@
The identity map and some graphs of $\operatorname{L}_{q}(A,B)$ for $q\in\mathbb{C}$ with $\lvert q\rvert<1+\epsilon$.